OpenStudy (anonymous):

Solve for x (5x+2)(3x-6)=0

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

How did you arrive at that answer?

OpenStudy (anonymous):

In these cases, multiply everything from one bracket to the other. 5x x 3x = ? 5x x -6 = ? 2 x 3x = ? 2 x -6 = ?

OpenStudy (anonymous):

0.4

OpenStudy (anonymous):

negative rather

OpenStudy (anonymous):

-0.4

OpenStudy (anonymous):

15x^2 -24x-12?

OpenStudy (anonymous):

15x-30x + 6x -12 = 0 -9x = 12 9x = 12 x =-0.75

OpenStudy (anonymous):

Sorryf orgot the negative sign on the 12

OpenStudy (anonymous):

I'll show you in drawing how it's done.

OpenStudy (anonymous):



OpenStudy (anonymous):

But 15x^2 -24x -12 is still equal to zero (forgot to put that in the earlier post) so 15x^2 -24x-12=0, I don't see what can be done when 15x^2 -24x=12

OpenStudy (anonymous):

separate the two 5x+2=0 5x = -2 x = -2/5 3x-6=0 3x=6 x=2

OpenStudy (anonymous):

What if I factor out a 3 from 15x^2 -24x-12=0 so that 3(5x^2 -8x-4)=0 ?